2x+3y=200
5x+4y+290

Answer by mananth(16946)

2 x + 3 y = 200 .............1

5 x + 4 y = 290 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 3
-8 x -12 y = -800
15 x + 12 y = 870
Add the two equations
7 x = 70
/ 7
x = 10
plug value of x in (1)
2 x + 3 y = 200
20 + 3 y = 200
3 y = 200 -20
3 y = 180
y = 60
x= 10
y= 60
m.ananth@hotmail.ca